OpenStudy (idealist10):

For what values of k0 and k1 does the initial value problem x^2*y"-4xy'+6y=0, y(0)=k0, y'(0)=k1 have a solution?

OpenStudy (idealist10):

@pooja195 @Michele_Laino @campbell_st @uri @timo86m

OpenStudy (idealist10):

@mathstudent55 @whpalmer4 @just_one_last_goodbye @sammixboo @shifuyanli

OpenStudy (idealist10):

@amistre64 @freckles @dan815 @KendrickLamar2014 @whpalmer4 @sammixboo

OpenStudy (amistre64):

does laplace work for this?

OpenStudy (amistre64):

dividing by x^2 would mean that x=0 is a sore subject ...

OpenStudy (amistre64):

x^2*y"-4xy'+6y=0 let y=x^n x^2*n(n-1)x^(n-2)-4x nx^(n-1)+6x^n=0 n(n-1)x^n-4 nx^n+6x^n=0 x^n [ n(n-1) -4n +6]=0 n^2-5n+6=0 (n-3)(n-2)=0, n=3 or 2 y=ax^3 + bx^2 is our general solution

OpenStudy (idealist10):

Now what? And then?

OpenStudy (amistre64):

determine of y(0) and y'(0) have a solution set ... of course.

OpenStudy (idealist10):

I got it! Thanks for the help!